1. Sam maximizes his utility by spending his entire income of $16 on bats and lizards.
Table below has Sam’s marginal utility from each good. The price of a bat is $2, and he
buys 2 bats. The marginal utility from the last lizard he buys is 36.
Quantity: Bats: Lizards:
MU: and TU: MU: and TU:
1 20 ? 50 ?
2 18 ? 44 ?
3 14 ? 38 ?
4 8 ? 26 ?
a. Fill in total utility for each good.

b. Calculate the price of a lizard. (Hint: use the concept of marginal utility per dollar)

c. What is Sam’s total utility if he maximizes his utility by spending his income of $16?

d. Sam could buy 4 lizards. If he did so and also purchased the maximum number of
bats possible given his income, what would be his total utility then?
1
Expert's answer
2013-03-29T09:49:01-0400
1. Sam maximizes his utility by spending his entire income of $16 on bats and lizards.Table below has Sam’s marginal utility
from each good. The price of a bat is $2, and he buys 2 bats. The
marginal utility from the last lizard he buys is 36.
Quantity: MU: and TU (Bats) MU: and TU (Lizards)
1 20
20 50
50
2 18
38 44
94
3 14
52 36
130
4 8
60 26
156
a. Fill in total utility for each good.

b. Calculate the price of a lizard. (Hint: use the concept of
marginal utility per dollar)

Concept of marginal utility per dollar: MUa/Pa = MUb/Pb
For2 Bats MU/P = 18/2$ = 9
MUbats/Pbats= MUlizards/Plizards
MUlizards/Plizards= 9
2*2$+ Plizards*Q = 16
Plizards*Q= 12
MUlizards/Plizards= 9
P= 4, Q = 3, MU = 36
c. Whatis Sam’s total utility if he maximizes his utility by spending his income
of $16?
TUtotal= 38 + 130 = 168

d. Sam could buy 4 lizards. If he did so and also purchased the
maximum number of bats possible given his income, what would be his
total utility then?
TUtotal= TU of 2 bats + TU of 4 lizards = 38 + 130 = 168